the following information about an economy is givicen: C=100+0.7YD, I=80-150i, G=60, T=40, Md=Y-3000i and Ms=1000. a. Derive the IS curve b. Derive the LM curve c. find the equilibruim value of Y and i


1
Expert's answer
2021-07-30T08:36:02-0400

a

"\text{Y=C+I+G}\\
\text{Y=100+0.7YD+80-150i+60}\\
\text{Y-0.7YD=240-150i}\\
Y=\frac{240-150i}{1-0.7D}\\"


b

"Y-3000i=1000\\
Y=1000+3000i\\"


c

"10000+250i=4250-125i"

"\text{i}=\dfrac{5750}{375}=15.33\%"

"\text{Y}=0.8(1-0.25)Y+900-50i+1000"

"\text{Y}=0.6Y+1900-50i"

"Y=4750-125i"

Equilibrium interest rate is:

"4750-125i=250i+10000"

"i=\dfrac{5250}{375}=14\%"


Equilibrium national income will, therefore, be:

"Y=4750-125\times 0.14=4732.5"